# The Composition and Chemistry of Titan’s Atmosphere 111Keywords: Titan; Astrochemistry; Astrobiology; Atmospheres; Photochemistry

### v1

@@ 1 Introduction
First and foremost, it is a moon with a dense atmosphere (Fig. 1), the only such body known in our Solar System. Also that this atmosphere, composed primarily of molecular nitrogen and methane, is a largely anoxic environment, with little oxygen to cause the termination of complex organic reactions. The result is a chemical wonderland, with a breathtaking array of complex organic molecules, of which we presently have only the most rudimentary understanding. Fig. 2 shows a schematic overview of the presumed chemistry that occurs in Titan’s upper atmosphere, where the ‘raw ingredients’ of its photochemical reactions, N2 and CH4, are broken apart and recombined into successively larger molecules, and finally haze particles.
@@
Achieving a better understanding of Titan’s atmosphere and its chemistry is important both for the sake of Titan science itself, and because of its potential to inform us about other environments. This includes the present-day Earth, since Titan and Earth are the only objects in the Solar System today to have a hydrological cycle of evaporation, condensation and precipitation, and associated rivers, lakes and seas 4, 5, 6, 7. Like the Earth, Titan also experiences seasons, due to orbiting close to Saturn’s equatorial plane, which is tilted 27∘ to the ecliptic. Titan therefore experiences summer and winter in each hemisphere, seasons that last 7.4 longer than on Earth, with transitional equinox periods of equal daylight at all latitudes.
@@ 2.1 Atmospheric Composition
Titan’s atmosphere is largely composed of two gases: N2 and CH4. The vertical profile of methane comes from measurements by instruments on the Cassini-Huygens space mission, primarily the Huygens GCMS (Gas Chromatograph and Mass Spectrometer) 24 from 0–146 km, the Cassini Visual and Infrared Mapping Spectrometer (VIMS)25 (50–850 km), the Cassini UVIS (Ultraviolet Imaging Spectrometer) 26 (400–1650 km), and the Cassini INMS (Ion and Neutral Mass Spectrometer) 27 (900–1500 km). Their results have been reported in publications from the mission 28, 29, 30, 31.
Aside from noble gases (36Ar, 40Ar and 22Ne) 32, 31, 22 molecular species other than N2 and CH4 have been definitively detected in Titan’s atmosphere at the time of writing (see Fig. 3): ten hydrocarbons (C2H2, C2H4, C2H6, c-C3H2, CH2CCH2, CH3CCH, C3H6, C3H8, C4H2, c-C6H6), eight cyanides222Also referred to as ‘nitriles’ when -CN occurs in an organic molecule. (HCN, HNC, HC3N, C2N2, CH3CN, C2H3CN, C2H5CN, CH3C3N), three oxygen-bearing species (CO, CO2, H2O) plus H2. These gases were originally detected by a variety of astronomical and remote sensing techniques from the ground and space.
@@
Oxygen has yet to be detected on Titan in an organic molecule such as methanol (CH3OH) or formaldehyde (H2CO), being found so far only in the simple inorganic molecules CO, CO2, and H2O. This limits the presently confirmed scope of astrobiological molecules (i.e. those with the elements CHON in a variety of functional groups) - at least in the atmosphere. At the surface and in the subsurface - where hydrocarbons are thought to be readily hydrolyzed as seen in laboratory experiments 33, 34, 35 - the astrobiological potential may be much greater 36, 37, 38.
@@ 2.2 Atmospheric Temperature Structure
Temperatures fall throughout the next layer, the mesosphere, as haze becomes thin, and radiative cooling by gases such as HCN and C2H2 becomes increasingly important 129, 135. Titan’s mesopause is reached at 600 km 128, above which altitude temperatures rise again. This is primarily due to methane UV absorption 138, blocking of outgoing IR radiation by C2H6, and far-infrared HCN rotational lines 129. This is the thermosphere, a region where gas collisions are rare, and molecules must wait to spontaneously emit a photon to lose energy.
@@ 2.3 Gas Vertical Profiles
Fig. 5 shows typical vertical profiles of the 24 known molecular species at low latitudes, compiled from a combination of ground and space-based measurements, and some photochemical model profiles constrained by observations. These include two pairs of structural isomers: HCN and HNC, CH3CCH and CH2CCH2.
Some important trends can be noted. Methane is an unusual outlier, with a greater abundance in the troposphere ( km) than above. This is due to the ‘cold trap’ effect, where it reaches saturation as the tropospheric temperature drops with altitude, and therefore its mixing ratio is reduced as it forms clouds at 15-30 km. The gases N2 and CO are well-mixed, having approximately uniform profiles throughout the atmosphere due to long photochemical lifetimes. Two other gases: H2 and C2H4 also do not condense at the ‘cold trap’ - the coldest part of the atmosphere around the tropopause at 45 km (70 K). The profile of H2 is shown here as constant at the 0.1% value typical of the lower stratosphere, since measurements of its vertical profile remains uncertain.
@@ 4 Chemistry of the Neutral Atmosphere
Note on chemical names: the topic of chemical nomenclature remains eternally problematic. For example, the relatively simple compound C2H3CN has been commonly referred to as ‘vinyl cyanide’ in most 20th century astronomical literature, though preference has shifted more recently towards the simpler, single-word name ‘acrylonitrile’ (also with ‘methyl cyanide’ to ‘acetonitrile’). In fact multiple valid names for C2H3CN exist, including ‘2-propenenitrile’, ‘cyanoethene/ cyanoethylene’ and ‘propenenitrile’, however it must be noted that the preferred official (IUPAC) name is actually the rather cumbersome ‘prop-2-enenitrile’.
@@ 4.1 Hydrocarbons and Hydrogen
Cyclic hydrocarbons occur where there are one or more closed rings of carbon atoms (at least three are needed to make a ring, or cycle). Cyclic molecules that have a pi bond of delocalized electrons are known as ‘aromatic’ (small aromatics are typically volatile at room temperature), although not all rings are aromatic. For example the unsaturated six-carbon ring benzene (c-C6H6) is a well-known aromatic, whereas its saturated cousin cyclohexane (c-C6H12) is a non-aromatic cycle. Carbenes are reactive molecules where carbon has two unbonded but paired electrons, such as methylene (CH2).
@@
leading to the trace amounts of hydrogen (0.1%) found in Titan’s lower atmosphere. Note that this termolecular reaction is an important process leading to the creation of H2 in the Interstellar Medium (ISM).
@@ 4.1.1 Hydrogen
Detection: Molecular hydrogen (H2) was was tentatively detected by Trafton233 using the 107 inch telescope at the McDonald Observatory, via absorption in the S(0) and S(1) quadrupole lines of the 3-0 band at 0.82 m. Hydrogen was clearly detected in the far-infrared by the Voyager 1 IRIS spectrometer 234 and confirmed by Cassini CIRS 235. The H2 Volume Mixing Ratio (VMR) was measured directly by the Cassini INMS instrument in the ionosphere as 0.4% 28 and in the lower stratosphere and troposphere by Huygens GCMS at 0.1% 31. The fact that hydrogen, presumed to be produced in the upper atmosphere by photolysis of methane 40, 41, was measured to have a decreasing abundance downwards, has proved difficult to replicate in models. Models have required a sink for H2 at the surface 225, 135, 236, which has even been suggested as possibly biological in origin 237. Important reactions for hydrogen are shown in Fig. 10 238.
@@ 4.1.2 Methane
Profile: Methane (CH4) was the first molecule to be positively identified in Titan’s atmosphere, via visible and near-IR absorptions seen by Gerard Kuiper 246, using the 82 inch reflector at McDonald Observatory. We now know that methane is the basic ingredient enabling all of Titan’s complex organic (i.e. carbon) chemistry, allowing reactions to proceed up to the creation of haze particles: some key reactions for methane are shown in Fig. 11247, 248.
@@
Loss: The photolysis of methane in the upper atmosphere leads to the formation of radicals including methyl (CH3), methylidene (CH), and the carbene methylene ( or )103 which undergo a chain of reactions to form all the hydrocarbons found in Titan’s atmosphere. Note that as much as 75% of methane photolysis above 700 km is due to the solar Lyman- line at 121.6 nm.257
@@ 4.1.3 Acetylene
Finally, neutral C2H2 in the ionosphere may combine with other ions to build heavier species239, e.g.:
@@
Acetylene absorbs photons to longer wavelengths (230 nm79) than methane, so its photolysis continues into the stratosphere (see Fig. 12). This produces ethynyl (C2H) and carbyne (C2) which are potent means of methane depletion via hydrogen abstraction:223
@@
C2H2 was one of the first molecules investigated to form a co-crystal 276, 277 at Titan surface temperatures, an organized co-condensate of two or more chemical species. The validity of multiple co-crystal types has since been established, but requires further laboratory work to determine the full parameter space of possible crystalline types.
@@ 4.1.4 Ethylene
Ethylene (C2H4) was discovered by infrared spectroscopy at the same time as acetylene 259. The vertical profile of ethylene exhibited a surprising trend to decrease in abundance upwards in the lower stratosphere early in the Cassini mission163, 278, although this faded at later seasons 137, 127.
Ethylene is a crucial, two-carbon neutral molecule that provides a stepping stone from methane to higher hydrocarbons (Fig. 13)279, 280. Ethylene is remarkable in being one of the few molecules (along with H2 and N2) that does not condense at the tropopause, and therefore persists in significant quantities into the troposphere.
@@ 4.1.6 Cyclopropenylidene
Cyclopropenylidene (c-C3H2) is the first carbene (a molecule having two unbonded, self-paired valence electrons from a carbon atom) and the second cyclic molecule to be found in Titan’s atmosphere (after benzene). c-C3H2 was detected using millimeter wavelength astronomy with ALMA 165, the third molecule whose first detection on Titan was achieved with with this telescope.
@@
which then recombines with to produce c-C3H2 306 and H.
@@
c-C3H2 can also result from collisional isomerization from its isomers propynylidene (t-C3H2) and propadienylidene (l-C3H2):311
which in turn may dissociatively recombine with an electron, returning to c-C3H2, or splinter into smaller acyclic fragments:266, 313
In the neutral atmosphere above 600 km, c-C3H2 photodissociates to form c-C3H, l-C3H, and C3. It is also lost by reaction with CH3 to form acylic species:311
Below 600 km, these pathways continue to be significant, but there is additional c-C3H2 loss via:311
Future work: Chemical pathways leading to and from c-C3H2 in Titan’s atmosphere remain to be explored, especially given the multiple possible structures for , including propynylidene ( ), propadienylidene (), cyclopropyne () and propenediylidene () 309, 310. Experimental and theoretical work on reaction pathways, rates and branching rates during photolysis will greatly help to clarify the production, loss and stability of cyclopropenylidene.
Several isomers of have been detected in space, including H2CCC 314, which should prompt further astronomical observations to determine if these isomers also exist in Titan’s upper atmosphere. Intriguingly, the CHCCH form has been shown to dimerize to form para-benzyne 307, discussed in a later section.
@@ 4.1.7 Propyne
The propyne (CH3CCH, methyl acetylene) isomer of () was first detected in the infrared by the Voyager Infrared Interferometer Spectrometer (IRIS)315 via long wavelength infrared emission bands at 328 and 633 cm-1316.
Production: Propyne and its symmetric isomer propadiene (CH2CCH2, allene) are produced by addition of CH into ethylene (Eq. 79, Fig. 16)317, 318, 319, but also through H-addition to propargyl320:
In the ionosphere, the C3H ion is a precursor to C3H4, and produced via:219
@@
Future work: Many reactions forming or depleting have uncertain branching ratios between CH3CCH and its isomer CH2CCH2. Further work is needed to improve knowledge of these quantities. Collisional interconversion between the two isomers may be mediated by atomic hydrogen99, so accurate measurement of both isomers may be a way to provide constraint on the abundance of otherwise short-lived and difficult to measure H atom.
@@ 4.1.8 Propadiene
Propadiene is a less abundant and less thermodynamically stable isomer of C3H4, which is more abundant in Titan’s atmosphere in the form of propyne. Propadiene (CH2CCH2) was detected in Titan’s atmosphere using high-resolution ground-based spectroscopy at NASA’s Infrared Telescope Facility (IRTF) with the Texas Echelon Cross Echelle Spectrograph (TEXES) instrument via its band at 845 cm-1 325.
Production: Like propyne (CH3CCH), propadiene is produced in the upper atmosphere by CH addition to ethylene (Eq. 79, Fig. 17)326, 318, 327, 328, and by H addition to C3H3:320
@@
Ion formation pathways of CH2CCH2 are less certain, but may follow similar channels as propyne, with branching ratios that are currently uncertain.
@@ 4.1.9 Propene
Propene (C3H6, propylene) was first detected using data from the Cassini CIRS infrared spectrometer via its band emission near 11 m 330, 171.
@@ 4.1.10 Propane
Propane (C3H8) was detected contemporaneously with propyne (CH3CCH) by Voyager’s IRIS instrument 316 via an infrared band at 748 cm-1 and subsequently confirmed by ground-based observations335 and with Cassini CIRS 286, 163, 336.

### v2

# The Composition and Chemistry of Titan’s Atmosphere 111Accepted for publication in ACS Earth and Space Chemistry

@@ 1 Introduction
First and foremost, it is a moon with a dense atmosphere (Fig. 1), the only such body known in our Solar System. Also that this atmosphere, composed primarily of molecular nitrogen and methane, is a largely anoxic environment, with little oxygen to cause the termination of complex organic reactions. The result is a chemical wonderland, with a breathtaking array of complex organic molecules, of which we presently have only the most rudimentary understanding. Fig. 2 shows a schematic overview of the presumed chemistry that occurs in Titan’s upper atmosphere, where the ‘raw ingredients’ of its photochemical reactions, N and CH, are broken apart and recombined into successively larger molecules, and finally haze particles.
@@
Achieving a better understanding of Titan’s atmosphere and its chemistry is important both for the sake of Titan science itself, and because of its potential to inform us about other environments. This includes the present-day Earth, since Titan and Earth are the only objects in the Solar System today to have a hydrological cycle of evaporation, condensation and precipitation, and associated rivers, lakes and seas 4, 5, 6, 7. Like the Earth, Titan also experiences seasons, due to orbiting close to Saturn’s equatorial plane, which is tilted 27 to the ecliptic. Titan therefore experiences summer and winter in each hemisphere, seasons that last 7.4 longer than on Earth, with transitional equinox periods of equal daylight at all latitudes.
@@ 2.1 Atmospheric Composition
Titan’s atmosphere is largely composed of two gases: N and CH. The vertical profile of methane comes from measurements by instruments on the Cassini-Huygens space mission, primarily the Huygens GCMS (Gas Chromatograph and Mass Spectrometer) 24 from 0–146 km, the Cassini Visual and Infrared Mapping Spectrometer (VIMS)25 (50–850 km), the Cassini UVIS (Ultraviolet Imaging Spectrometer) 26 (400–1650 km), and the Cassini INMS (Ion and Neutral Mass Spectrometer) 27 (900–1500 km). Their results have been reported in publications from the mission 28, 29, 30, 31.
Aside from noble gases (Ar, Ar and Ne) 32, 31, 22 molecular species other than N and CH have been definitively detected in Titan’s atmosphere at the time of writing (see Fig. 3): ten hydrocarbons (CH, CH, CH, c-CH, CHCCH, CHCCH, CH, CH, CH, c-CH), eight cyanides222Also referred to as ‘nitriles’ when -CN occurs in an organic molecule. (HCN, HNC, HCN, CN, CHCN, CHCN, CHCN, CHCN), three oxygen-bearing species (CO, CO, HO) plus H. These gases were originally detected by a variety of astronomical and remote sensing techniques from the ground and space.
@@
Oxygen has yet to be detected on Titan in an organic molecule such as methanol (CHOH) or formaldehyde (HCO), being found so far only in the simple inorganic molecules CO, CO, and HO. This limits the presently confirmed scope of astrobiological molecules (i.e. those with the elements CHON in a variety of functional groups) - at least in the atmosphere. At the surface and in the subsurface - where hydrocarbons are thought to be readily hydrolyzed as seen in laboratory experiments 33, 34, 35 - the astrobiological potential may be much greater 36, 37, 38.
@@ 2.2 Atmospheric Temperature Structure
Temperatures fall throughout the next layer, the mesosphere, as haze becomes thin, and radiative cooling by gases such as HCN and CH becomes increasingly important 129, 135. Titan’s mesopause is reached at 600 km 128, above which altitude temperatures rise again. This is primarily due to methane UV absorption 138, blocking of outgoing IR radiation by CH, and far-infrared HCN rotational lines 129. This is the thermosphere, a region where gas collisions are rare, and molecules must wait to spontaneously emit a photon to lose energy.
@@ 2.3 Gas Vertical Profiles
Fig. 5 shows typical vertical profiles of the 24 known molecular species at low latitudes, compiled from a combination of ground and space-based measurements, and some photochemical model profiles constrained by observations. These include two pairs of structural isomers: HCN and HNC, CHCCH and CHCCH.
Some important trends can be noted. Methane is an unusual outlier, with a greater abundance in the troposphere ( km) than above. This is due to the ‘cold trap’ effect, where it reaches saturation as the tropospheric temperature drops with altitude, and therefore its mixing ratio is reduced as it forms clouds at 15-30 km. The gases N and CO are well-mixed, having approximately uniform profiles throughout the atmosphere due to long photochemical lifetimes. Two other gases: H and CH also do not condense at the ‘cold trap’ - the coldest part of the atmosphere around the tropopause at 45 km (70 K). The profile of H is shown here as constant at the 0.1% value typical of the lower stratosphere, since measurements of its vertical profile remains uncertain.
@@ 4 Chemistry of the Neutral Atmosphere
Note on chemical names: the topic of chemical nomenclature remains eternally problematic. For example, the relatively simple compound CHCN has been commonly referred to as ‘vinyl cyanide’ in most 20th century astronomical literature, though preference has shifted more recently towards the simpler, single-word name ‘acrylonitrile’ (also with ‘methyl cyanide’ to ‘acetonitrile’). In fact multiple valid names for CHCN exist, including ‘2-propenenitrile’, ‘cyanoethene/ cyanoethylene’ and ‘propenenitrile’, however it must be noted that the preferred official (IUPAC) name is actually the rather cumbersome ‘prop-2-enenitrile’.
@@ 4.1 Hydrocarbons and Hydrogen
Cyclic hydrocarbons occur where there are one or more closed rings of carbon atoms (at least three are needed to make a ring, or cycle). Cyclic molecules that have a pi bond of delocalized electrons are known as ‘aromatic’ (small aromatics are typically volatile at room temperature), although not all rings are aromatic. For example the unsaturated six-carbon ring benzene (c-CH) is a well-known aromatic, whereas its saturated cousin cyclohexane (c-CH) is a non-aromatic cycle. Carbenes are reactive molecules where carbon has two unbonded but paired electrons, such as methylene (CH).
@@
leading to the trace amounts of hydrogen (0.1%) found in Titan’s lower atmosphere. Note that this termolecular reaction is an important process leading to the creation of H in the Interstellar Medium (ISM).
@@ 4.1.1 Hydrogen
Detection: Molecular hydrogen (H) was was tentatively detected by Trafton233 using the 107 inch telescope at the McDonald Observatory, via absorption in the S(0) and S(1) quadrupole lines of the 3-0 band at 0.82 m. Hydrogen was clearly detected in the far-infrared by the Voyager 1 IRIS spectrometer 234 and confirmed by Cassini CIRS 235. The H Volume Mixing Ratio (VMR) was measured directly by the Cassini INMS instrument in the ionosphere as 0.4% 28 and in the lower stratosphere and troposphere by Huygens GCMS at 0.1% 31. The fact that hydrogen, presumed to be produced in the upper atmosphere by photolysis of methane 40, 41, was measured to have a decreasing abundance downwards, has proved difficult to replicate in models. Models have required a sink for H at the surface 225, 135, 236, which has even been suggested as possibly biological in origin 237. Important reactions for hydrogen are shown in Fig. 10 238.
@@ 4.1.2 Methane
Profile: Methane (CH) was the first molecule to be positively identified in Titan’s atmosphere, via visible and near-IR absorptions seen by Gerard Kuiper 246, using the 82 inch reflector at McDonald Observatory. We now know that methane is the basic ingredient enabling all of Titan’s complex organic (i.e. carbon) chemistry, allowing reactions to proceed up to the creation of haze particles: some key reactions for methane are shown in Fig. 11247, 248.
@@
Loss: The photolysis of methane in the upper atmosphere leads to the formation of radicals including methyl (CH), methylidene (CH), and the carbene methylene ( or )103 which undergo a chain of reactions to form all the hydrocarbons found in Titan’s atmosphere. Note that as much as 75% of methane photolysis above 700 km is due to the solar Lyman- line at 121.6 nm.257
@@ 4.1.3 Acetylene
Finally, neutral CH in the ionosphere may combine with other ions to build heavier species239, e.g.:
@@
Acetylene absorbs photons to longer wavelengths (230 nm79) than methane, so its photolysis continues into the stratosphere (see Fig. 12). This produces ethynyl (CH) and carbyne (C) which are potent means of methane depletion via hydrogen abstraction:223
@@
CH was one of the first molecules investigated to form a co-crystal 276, 277 at Titan surface temperatures, an organized co-condensate of two or more chemical species. The validity of multiple co-crystal types has since been established, but requires further laboratory work to determine the full parameter space of possible crystalline types.
@@ 4.1.4 Ethylene
Ethylene (CH) was discovered by infrared spectroscopy at the same time as acetylene 259. The vertical profile of ethylene exhibited a surprising trend to decrease in abundance upwards in the lower stratosphere early in the Cassini mission163, 278, although this faded at later seasons 137, 127.
Ethylene is a crucial, two-carbon neutral molecule that provides a stepping stone from methane to higher hydrocarbons (Fig. 13)279, 280. Ethylene is remarkable in being one of the few molecules (along with H and N) that does not condense at the tropopause, and therefore persists in significant quantities into the troposphere.
@@ 4.1.6 Cyclopropenylidene
Cyclopropenylidene (c-CH) is the first carbene (a molecule having two unbonded, self-paired valence electrons from a carbon atom) and the second cyclic molecule to be found in Titan’s atmosphere (after benzene). c-CH was detected using millimeter wavelength astronomy with ALMA 165, the third molecule whose first detection on Titan was achieved with with this telescope.
@@
which then recombines with to produce c-CH 306 and H.
@@
c-CH can also result from collisional isomerization from its isomers propynylidene (t-CH) and propadienylidene (l-CH):311
which in turn may dissociatively recombine with an electron, returning to c-CH, or splinter into smaller acyclic fragments:266, 313
In the neutral atmosphere above 600 km, c-CH photodissociates to form c-CH, l-CH, and C. It is also lost by reaction with CH to form acylic species:311
Below 600 km, these pathways continue to be significant, but there is additional c-CH loss via:311
Future work: Chemical pathways leading to and from c-CH in Titan’s atmosphere remain to be explored, especially given the multiple possible structures for , including propynylidene ( ), propadienylidene (), cyclopropyne () and propenediylidene () 309, 310. Experimental and theoretical work on reaction pathways, rates and branching rates during photolysis will greatly help to clarify the production, loss and stability of cyclopropenylidene.
Several isomers of have been detected in space, including HCCC 314, which should prompt further astronomical observations to determine if these isomers also exist in Titan’s upper atmosphere. Intriguingly, the CHCCH form has been shown to dimerize to form para-benzyne 307, discussed in a later section.
@@ 4.1.7 Propyne
The propyne (CHCCH, methyl acetylene) isomer of () was first detected in the infrared by the Voyager Infrared Interferometer Spectrometer (IRIS)315 via long wavelength infrared emission bands at 328 and 633 cm316.
Production: Propyne and its symmetric isomer propadiene (CHCCH, allene) are produced by addition of CH into ethylene (Eq. 79, Fig. 16)317, 318, 319, but also through H-addition to propargyl320:
In the ionosphere, the CH ion is a precursor to CH, and produced via:219
@@
Future work: Many reactions forming or depleting have uncertain branching ratios between CHCCH and its isomer CHCCH. Further work is needed to improve knowledge of these quantities. Collisional interconversion between the two isomers may be mediated by atomic hydrogen99, so accurate measurement of both isomers may be a way to provide constraint on the abundance of otherwise short-lived and difficult to measure H atom.
@@ 4.1.8 Propadiene
Propadiene is a less abundant and less thermodynamically stable isomer of CH, which is more abundant in Titan’s atmosphere in the form of propyne. Propadiene (CHCCH) was detected in Titan’s atmosphere using high-resolution ground-based spectroscopy at NASA’s Infrared Telescope Facility (IRTF) with the Texas Echelon Cross Echelle Spectrograph (TEXES) instrument via its band at 845 cm 325.
Production: Like propyne (CHCCH), propadiene is produced in the upper atmosphere by CH addition to ethylene (Eq. 79, Fig. 17)326, 318, 327, 328, and by H addition to CH:320
@@
Ion formation pathways of CHCCH are less certain, but may follow similar channels as propyne, with branching ratios that are currently uncertain.
@@ 4.1.9 Propene
Propene (CH, propylene) was first detected using data from the Cassini CIRS infrared spectrometer via its band emission near 11 m 330, 171.
@@ 4.1.10 Propane
Propane (CH) was detected contemporaneously with propyne (CHCCH) by Voyager’s IRIS instrument 316 via an infrared band at 748 cm and subsequently confirmed by ground-based observations335 and with Cassini CIRS 286, 163, 336.